### v4

# PoI: A Filter to Extract Pixel of Interest from Novel View Synthesis for Scene Coordinate Regression

Abstract
Novel View synthesis (NVS) techniques, notably Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS), can augment camera pose estimation by extending training data with rendered images. However, the images rendered by these methods are often plagued by blurring, undermining their reliability as training data for camera pose estimation. This limitation is particularly critical for Scene Coordinate Regression (SCR) methods, which aim at pixel-level 3D coordinate estimation, because rendering artifacts directly lead to estimation inaccuracies. To address this challenge, we propose a dual-criteria filtering mechanism that dynamically identifies and discards suboptimal pixels during training. The dual-criteria filter evaluates two concurrent metrics: (1) real-time SCR reprojection error, and (2) gradient threshold, across the coordinate regression domain. In addition, for visual localization problems in sparse input scenarios, it will be even more necessary to use data generated by NVS to assist the localization task. We design a coarse-to-fine PoI variant using sparse input NVS to solve this problem. Experiments across indoor and outdoor benchmarks confirm our method’s efficacy. It achieves state-of-the-art localization accuracy while maintaining computational efficiency.
I Introduction
Figure 1: Left: Comparison of query and rendered images of the dataset 7Scenes and Cambridge Landmarks, revealing uneven quality within them, with query frames clear and rendered frames blurry; Right: Translation error versus training time, where ‘CoodiNet+’ denotes using rendered images as query images for CPR method CoodiNet (LENS in this case); ‘DSAC*+’ and ‘ACE+’ denote the method that combines NVS-rendered images and query images as training data for SCR method DSAC* and ACE. ‘ACE+PoI’ denotes our proposed PoI method (ACE-based); Analysis reveals that directly adding rendered data to the training set will increase training time to some extent, but performance will decrease for the SCR method. On the other hand, our PoI approach can improve the performance with an acceptable time increase.
Visual localization, or camera pose estimation, is a fundamental task in computer vision that involves estimating the 6-degree-of-freedom (6DoF) camera poses within a known scene based on input images. This task plays a crucial role in Simultaneous Localization and Mapping (SLAM) [1, 2, 3, 4, 5] and has wide applications in areas such as autonomous driving, robotics, and virtual reality.
Learning-based methods for visual localization can be categorized into two main types: Camera Pose Regression (CPR) methods [6, 7, 8, 9, 10, 11, 12], which directly regress the camera pose from the input image, and Scene Coordinate Regression (SCR) methods [13, 14, 15, 16, 17, 18], which predict the 3D coordinates of the pixels of the images and then estimate the camera pose based on the 2D-3D correspondence. However, both methods require the training data to have a suitable sampling distribution and sufficient sampling density to ensure reliable pose estimation of test images. It is even more difficult to obtain the corresponding camera pose labels.
In light of these requirements, some CPR-based methods try to enrich the training dataset with synthetic data rendered by NVS techniques, which are referred to as CPR-based Neural Rendering Pose (NRP) estimation methods. For example, LENS [11] employs NeRF to render views of uniformly sampled novel poses, thereby augmenting the training dataset and treating these synthetic images as real data without additional processing. Similarly, DFNet [12] utilizes NeRF-W [19] for NVS and features a cross-domain design that helps to minimize the discrepancies between synthetic and query images, effectively bridging the gap between the two domains.
Although artifacts exist in rendered images, as shown in the left part of Figure 1, the CPR-based NRP methods achieve effectiveness mainly because they involve N-to-1 prediction (1 prediction for N pixels) and are more affected by the overall quality of the images than particular pixels. In contrast, SCR-based methods rely on accurate N-to-N prediction (N predictions for N pixels) and are particularly susceptible to the quality of rendered pixels. As shown in the right part of Figure 1, after extending the training dataset with synthetic data through NVS, the CPR method shows significant improvement, while the performance of SCR declines, accompanied by a notable increase in training time. Directly training the SCR model with raw rendered images proves less effective than CPR methods. As a result, very few studies can effectively combine NVS with the SCR framework to expand the training dataset.
To address the issue of SCR-based NRP, we designed a novel pixel of interest (PoI) module that acts as a run-time filter for the pixels from rendered images. Specifically, the 3D-to-2D projection error of each pixel along with the network gradient is employed as a criterion for whether the point is retained. As the training progresses, the PoI method gradually removes poorly rendered pixels and leverages the remaining points alongside real data to train the network.
Moreover, we propose a coarse-to-fine variant of the PoI method to address visual localization in scenarios of extremely sparse input. In the coarse stage, we receive all available synthetic data as input and gradually train a coarse model by self-pruning to remove noisy pixels. Following this, we fine-tune the coarse model using sparse real pixels. This enables our PoI variant method to efficiently leverage sparse input while ensuring strong pose estimation performance, even under difficult conditions.
The main contributions of our work are summarized as follows:
  • We introduce the PoI method, a pixel-level filter designed to eliminate pixels of low rendering quality for effective training data augmentation.
  • We present an innovative variant of the PoI method to tackle scene coordinate regression for scenarios of extremely sparse input.
  • We perform extensive evaluations on both indoor and outdoor datasets, demonstrating the effectiveness of our method and achieving state-of-the-art performance.
II Related Work
Camera Pose Regression The CPR methods, i.e., to regress the camera pose from the given image directly, are the most naive ideas in learning-based methods [20, 21, 22, 23, 24, 25, 26, 27, 12, 28]. The most straightforward method implicitly uses CNN layers or MLP to represent the image-to-pose correspondence. PoseNet [20] first proposes CPR using pre-trained GoogLeNet as the feature extractor. Then, several works focus on improving CPR through additional modules. Geomapnet [22] estimates the absolute camera poses and the relative poses between adjacent frames. AtLoc [25] uses a self-attention module to extract salient features from the image. Vlocnet++ [24] adds a semantic module to solve the dynamic scene and improve the robustness for blockings and blurs. Marepo [29] first regresses the scene-specific geometry from the input images and then estimates the camera pose using a scene-agnostic transformer. The CPR method has achieved excellent efficiency and framework simplification, but there is still room for improvement in accuracy. Scene Coordinate Regression The SCR-based method [16, 14, 30, 15, 31, 32, 13, 33] aims to estimate the camera pose in two steps: first, estimate the coordinates of the points in the 3D scene corresponding to the 2D pixels, and then estimate the camera pose based on the 2D-3D correspondence relationship. SCR was initially proposed using the random forest for RGB-D images [16]. Recently, estimating scene coordinates through RGB input has been widely studied. ForestNet[31] compares the benefits of Random Forest (RF) and Neural Networks in evaluating the scene coordinate and camera poses. ForestNet also proposes a novel method to initiate the neural network from an RF. DSAC [14], DSAC++ [30] devise a differentiable RANSAC, and thus the SCR method can be trained in an end-to-end manner. ESAC [15] uses a mixture of expert models (i.e., a gating network) to decide which domain the query belongs to, and then the complex SCR task can be split into simpler ones. DSAC* [13] extends previous works to applications using RGB-D or RGB images, with or without 3D models. This means that in the minimal case, only RGB images will be used as the input to DSAC*, just like most CPR methods. More information about the 3D structure will be utilized for most SCR methods than for CPR ones. However, approaches like DSAC* can achieve more accurate estimations even if the input is the same as the CPR method. ACE [18] and GLACE [34] abandon the time-consuming end-to-end supervision module and shuffle all pixels of the scene to improve training efficiency. ACE and GLACE utilize only RGBs without additional 3D geometry information and achieve comparable accuracy compared to former methods.
Neural Render Pose Estimation In many cases, collecting more high-quality and accurately annotated training data to improve localization accuracy is an effective solution. To reduce the cost and improve efficiency, Some works also try to use more flexible NVS to render synthetic images instead of collecting extra data [7, 8, 6, 9, 11, 12], where NVS is a set of techniques to render synthetic views from existing views. INeRF [35] applies an inverted NeRF to optimize the estimated pose through the color residual between rendered and observed images. However, the initially estimated poses are crucial in guaranteeing the convergence of outputs. LENS [11] samples the poses uniformly throughout the area and trains a NeRF-W [19] to render the synthetic images. Then, the rendered images and poses work as additional training data for the pose regression network. The limitation of LENS lies in the costly offline computation for dense samples. DFNet [12] uses direct feature matching between observed and synthetic images generated by histogram-assisted NeRF. The feature match approach is proposed to extract observed or generated images’ cross-domain information. All these methods combine the NVS module and the CPR module to optimize the performance of the absolute pose estimation of the photos.
However, since N-to-N prediction of SCR requires higher image fidelity than N-to-1 prediction of CPR, there is a lack of methods to effectively utilize the NVS-generated data in SCR-based methods. In this paper, we propose to use NVS-generated data in end-to-end SCR approaches such as ACE and GLACE. First, we design novel pose sampling methods to meet the multiple requirements of different datasets. To address the problem of varying lighting conditions, we use exposure histogram-assisted 3DGS as the baseline to sample new views of multiple exposures for each sampled pose in outdoor datasets. Second, we propose a pixel filter to remove bad pixels in rendered images and use query images and remaining rendered pixels to improve the estimation.
Figure 2: Pipeline of our proposed methods: (a) Data Prepare: We first sample a group of synthesized camera poses $P_{novel}$ according to the query training pose $P_{query}$ using Fisher Sample. Then, we render the synthesized views $I_{novel}$ based on the sampled poses $P_{novel}$ through the novel view synthesis model. (b) Architecture of PoI module: First, a pre-trained scene-irrelevant backbone is applied to extract the features of the input query photos $I_{query}$ and the synthesized novel images $I_{novel}$. Then, the filter is applied to the features of the rendered images and extracts the features of interest. After that, we combine the query features with the retained features of the novel views and shuffle the pixel-aligned features to get the aggregation. Finally, we estimate the scene coordinates of the pixels using a scene-specific Head. The filtering algorithm is designed based on the reprojection error and the gradient of the estimated scene coordinates.
III Method
In general, the pipeline of our proposed method is shown in Figure 2. For the input query images $I_{query}$ and corresponding camera poses $P_{query}$, we first sample the novel camera pose $P_{novel}$ using Fisher Sampling. Then we render novel views $I_{novel}$ through 3DGS. Finally, we use the PoI module to estimate the scene coordinates according to the input $I_{query},I_{novel}$. During the evaluation stage, we use PNP-based RANSAC to infer the camera poses from the scene coordinates.
The following part of this section is arranged as follows: We first introduce the building blocks of this paper, including NVS models, the synthetic pose sampling approach, and the baseline of the SCR regressor in Section III-A; Then we explain the details of the Fisher sample method in Section LABEL:method:fisher_sample; We elaborate on the architecture of the PoI method and introduce the details of the filtering strategy used in the PoI method in Section III-B and Section III-C; We finally explain the variant of the PoI method in extreme scenarios of sparse input in Section III-D.
III-A Building Blocks
Fisher Sample Method Before PoI training, we would sample novel camera poses $P_{novel}$ and render the corresponding images $I_{novel}$ based on the query images $I_{query}$ and the corresponding camera poses $P_{query}$. Since rendered novel views are expected to contain more information about the scenes beyond the training data, we choose to use the Fisher information selection method from FisherRF [36].
NVS: This paper uses 3DGS [37] as the baseline for novel view synthesis. A major challenge for 3DGS in outdoor scenarios is the failure caused by changes in lighting conditions, which also exists in the localization problem. To tackle this problem, we apply the luminance histogram method from DFNet [12] to adjust the appearance of the rendering images in 3DGS. We generate the exposure embedding from the luminance histogram and output the affine transformation from an MLP used in wild Gaussians [38].
Backbone: To expedite convergence, we use the pseudo-depth supervision mechanism from ACE for suboptimal pixels exhibiting elevated reprojection errors. Specifically, when a pixel’s reprojection error surpasses a predefined threshold, we initially enforce its convergence toward designated pseudo-depth targets through constrained optimization. After a certain epoch, when the estimated values of these pixels fall within the valid threshold, we use conventional supervision for further optimization.
III-B Architecture of PoI
Current approaches employing rendered images as auxiliary inputs for camera pose estimation indiscriminately process entire images, neglecting per-pixel rendering quality variations. This oversight induces dual inefficiencies: (1) Redundant low-fidelity pixel processing will incur significant computational overhead, and (2) Indiscriminate adoption of generated images will reduce the robustness of training. To address these limitations, we propose a discriminative filtering mechanism that selectively retains high-confidence rendered pixels. Our solution stems from the analysis of the reconstruction properties of 3D Gaussian distribution (3DGS): the lack of inter-pixel dependencies in RGB predictions leads to spatially uncorrelated rendering quality distributions across image planes. As a result, traditional frame-level filtering can lead to a catastrophic loss of high-fidelity pixels due to the elimination of entire images. Our methodological innovation establishes a per-pixel filtering paradigm through multi-frame quality consistency analysis. This method effectively decouples the pixel selection decision from the integrity constraints of a single frame.
The architectural framework of PoI is illustrated in Figure 2(b). The query image $I_{query}$ and rendered novel view $I_{novel}$ are processed through the backbone to yield corresponding query view features $F_{query}$ and novel view features $F_{novel}$. Subsequently, the proposed filter is implemented to obtain features of interest (FoI) within $F_{novel}$. These FoI descriptors are then fused with $F_{query}$ through feature concatenation, enabling precise regression of scene coordinates for the corresponding pixels via a coordinate prediction head.
III-C Filtering Strategy in PoI
The filtering pipeline operates through two coordinated mechanisms: Subsampling and Gradient-Reprojection Co-Filter method.
After data preparation, the amount of rendering data is the same as that of query data. During training, we implement the two-step progressive synthetic pixels filtering method to prevent model destabilization from imperfect renderings. First, our subsampling protocol extracts features from shuffled synthetic data with Bernoulli probability $p=0.5$.
After that, per-iteration filtration employs a dual-criterion gate function:
$\mathcal{G}(x,y)=\mathbb{I}\left[\|\nabla\theta\|_{2}<\tau_{g}\right]\odot\mathbb{I}\left[\mathcal{L}_{\text{reproj}}<\tau_{r}\right]$ (1)
where $\tau_{g},\tau_{r}$ are the pre-set thresholds of gradient magnitude and projection distance.
Because of the pseudo-depth introduced in the building blocks, the network gradient alone cannot represent real-time rendering quality. We also use reprojection error (the distance between pixel coordinates GT and the re-projected 2D pixel coordinates estimation) along with the gradient for filtering. Therefore, pixels with smaller gradients and lower reprojection loss are retained.
The filter will remove novel views features of outlier prediction. The remaining features are FoI, and the corresponding pixels of FoI are PoI. Figure 3 shows an example of the PoI results in the 7Scenes and Cambridge Landmarks datasets. After filtering, we combine and shuffle the features $F$ and FoI and put them into the scene-specific MLP Head to estimate the scene coordinates. Each pixel is saved with its corresponding frame ID and pixel coordinates to ensure realignment after shuffling.
It is worth mentioning that we have set a dynamic weight for the loss of rendered pixels. Because in the early stage of training, we want the model to converge quickly. After determining the PoI, we gradually reduce the weight of the loss of PoI from 1 to 0.01, while for the pixels from query images, we set the weight to 1 during the entire training process. The loss function is designed as follows:
$\begin{split}\mathcal{L}&=\begin{cases}\mathcal{L}_{rep}^{query}(i),&\text{if }i\in T\\ \tilde{\omega}\times\mathcal{L}_{rep}^{poi}(i),&\text{if }i\in PoI\end{cases}\\ \tilde{\omega}&=\omega_{max}-\frac{I_{iter}}{N_{iter}}(\omega_{max}-\omega_{min})\end{split}$ (2)
where $\mathcal{L}_{rep}$ means the reprojection loss, T denotes query data, $\tilde{\omega}$ denotes the dynamic weight of PoI loss, changing from $\omega_{max}$ (set to 1) to $\omega_{min}$ (set to 0.01). $I_{iter}$ denotes the current iteration number and $N_{iter}$ is the total iterations. All rendered pixels are initially set as PoI. As the training progresses, we rule out outlier prediction points from PoI. At the end of the training, the choice of PoI and the loss weight of PoI are fixed.
Figure 3: An example of the results of PoI method in dataset 7Scenes and Cambridge Landmarks. To highlight the determined pixels of interest, we scale up the ‘Value’ (V) of the HSV representation of the images.
Figure 4: Pipeline of the coarse-to-fine strategy for sparse input localization.
III-D Coarse-to-Fine Strategy for Extreme Sparse Input
Sparse-input visual localization (as few as 4-5 frames for some indoor scenes) presents significant challenges, as CPR and SCR techniques exhibit limitations in estimating unobserved scene regions. Their training paradigm is mainly based on 2D RGB input and lacks geometric constraints. To address this fundamental limitation, we propose leveraging sparse-input NVS to strategically extend the training views.
The primary technical challenge resides in the compromised visual quality of sparse-view NVS outputs compared to their dense-view counterparts. Direct application of the original Point of Interest (PoI) framework would disproportionately weight the implicit neural mapping process toward synthetic renderings, potentially compromising localization accuracy. Our solution introduces a novel coarse-to-fine training paradigm (illustrated in Figure 4) that systematically integrates synthetic and real data.
The initial coarse phase adopts the PoI module while exclusively utilizing the rendered images. A critical innovation in this stage is the implementation of a self-pruning mechanism with an elevated error tolerance threshold for reprojection loss filtering. The subsequent refinement stage employs curriculum learning to integrate query data with high-confidence synthetic samples progressively. Notably, we strategically maintain a reduced learning rate during fine-tuning compared with the coarse stage to enable detail-oriented optimization. This dual-phase approach eliminates filtering requirements in final training iterations, permitting full data utilization.
Refer to caption Refer to caption
(a) Coarse stage results. (b) Fine stage results.
Figure 5: The localization trajectories of the coarse-to-fine method for sparse view circumstances.
We construct the mesh based on the estimated scene coordinates of the coarse stage and the fine stage and visualize the camera pose estimation results in Figure 5. We may find that in the coarse stage, the pose estimation error is relatively large, and the quality of the reconstructed details is low. In the refined stage, the performance is much better.
Our framework demonstrates robust performance through comprehensive experimental validation across both indoor and outdoor benchmarking datasets. Quantitative evaluations and implementation specifics, including ablation studies on threshold selection and learning rate scheduling, are detailed in Section IV-D.
IV Experiment
TABLE I: Median errors of camera pose regression methods and scene coordinate regression methods on the 7Scenes dataset [16]. We bold the best results.
Method Scenes Avg. (cm/°)
Chess Fire Heads Office Pumpkin Kitchen Stairs
CPR PoseNet15 10/4.02 27/10.0 18/13.0 17/5.97 19/4.67 22/5.91 35/10.5 21/7.74
Marepo 1.9/0.83 2.3/0.92 2,1/1.24 2.9/0.93 2.5/0.88 2.9/0.98 5.9/1.48 2.9/1.04
SCR DSAC* 0.5/0.17 0.8/0.28 0.5/0.34 1.2/0.34 1.2/0.28 0.7/0.21 2.7/0.78 1.1/0.34
ACE 0.5/0.18 0.8/0.33 0.5/0.33 1.0/0.29 1.0/0.22 0.8/0.2 2.9/0.81 1.1/0.34
GLACE 0.6/0.18 0.9/0.34 0.6/0.34 1.1/0.29 0.9/0.23 0.8/0.20 3.2/0.93 1.2/0.36
NRP LENS 3/1.30 10/3.70 7/5.80 7/1.90 8/2.20 9/2.20 14/3.60 8/3.00
DFNet 3/1.12 6/2.30 4/2.29 6/1.54 7/1.92 7/1.74 12/2.63 6/1.93
GSplatLoc 0.43/0.16 1.03/0.32 1.06/0.62 1.85/0.4 1.8/0.35 2.71/0.55 8.83/2.34 2.53/0.68
PoI(ours) 0.5/0.13 0.7/0.31 0.5/0.25 0.9/0.27 1.0/0.16 0.7/0.21 2.4/0.68 1.0/0.28
GLPoI(ours) 0.5/0.10 0.8/0.28 0.4/0.27 0.8/0.24 0.8/0.22 0.6/0.20 2.5/0.77 0.9/0.29
IV-A Experiments Setup
TABLE II: Results on Cambridge Landmarks, because of the obvious gap between SCR-based methods and CPR-based methods, we only list SCR-based methods. column ‘Mapping time’ shows the training time of these methods (the GS model can be trained offline, and the training time of the GS model is shorter than PoI method), and column ‘Mapping size’ is the memory consumption for saving the parameters of the network. We bold the best result for group ‘SCR’ and group ‘NRP’ separately.
Method Mapping with Depth/Mesh Mapping Time Map Size Scenes Avg. (cm/°)
King’s Hospital Shop Church Court
SCR SANet [39] Yes  1min  260M 32/0.5 32/0.5 10/0.5 16/0.6 328/2 84/0.8
SRC [40] Yes 2min 40M 39/0.7 38/0.5 19/1 31/1.0 81/0.5 42/0.7
DSAC* [13] No 15h 28M 18/0.3 21/0.4 5/0.3 15/0.6 34/0.2 19/0.4
Poker [18] No 20min 16M 18/0.3 25/0.5 5/0.3 9/0.3 28/0.1 17/0.3
EGFS [33] No 21min 9M 14/0.3 28/0.1 19/0.4 5/0.2 10/0.3 15/0.3
GLACE [34] No 20min 13M 19/0.3 17/0.4 4/0.2 9/0.3 19/0.1 14/0.3
NRP LENS [11] No - - 33/0.5 44/0.9 27/1.6 53/1.6 - -
GSplatLoc [41] No - - 27/0.46 20/0.71 5/0.36 16/0.61 - -
PoI (ours) No 25min 16M 17/0.3 23/0.5 5/0.2 8/0.3 24/0.1 15/0.3
GLPoI (ours) No 25min 13M 18/0.3 15/0.4 4/0.2 8/0.3 17/0.1 12/0.3
The proposed framework is evaluated on two publicly available benchmarks: the Microsoft 7Scenes dataset [16] (with SFM-derived pseudo ground truth) and the Cambridge Landmarks dataset [20]. The methodology operates exclusively on RGB inputs paired with camera extrinsic parameters, excluding depth information from 7Scenes and structural reconstruction priors from Cambridge Landmarks. Full-resolution RGB images are preserved to ensure precise pose estimation capabilities.
To optimize computational efficiency, pose sampling and novel view synthesis are preprocessed offline, with generated camera poses and rendered images stored on disk. These precomputed rendered images are loaded with the original training data during the training phase. A specialized strategy is implemented for the ‘Kitchen’ scene, where training samples are partitioned into two clusters based on camera pose distributions. Following the ACE-inspired protocol, two distinct models are trained on these clusters. At inference time, the final pose estimation is selected from the model demonstrating superior RANSAC consensus through a higher proportion of inlier correspondences. The filter parameters $\tau_{g}$ and $\tau_{r}$ were set to 0.0005 and 5, respectively, for all scenes. All experiments utilized NVIDIA V100 GPUs. PoI training employed a single GPU, whereas GLPoI training used four GPUs in a distributed data-parallel setup to expedite convergence.
IV-B Quantitative Results
The comparison of median translation and rotation errors between our proposed methods with different camera pose regression methods (top), scene coordinate regression methods (middle), and neural render pose estimation methods (bottom) in dataset 7Scenes is shown in Table I. Generally speaking, scene coordinate regression methods outperform absolute camera pose regression methods in both translations and orientations. Unlike [41], we include DFNet and LENS in NRP, although they use offline NVS. Our proposed method outperforms DSAC* and ACE by exploiting the additional information from the rendered novel views. Our ‘GLPoI’ beats ‘GLACE’ and achieves the state of the art.
The results of the Cambridge Landmarks datasets are shown in Table II. Since apparent gaps exist between scene coordinate regression methods and absolute camera pose regression methods, we only list the results of SCR and NRP methods. We come to a similar conclusion as that of 7Scenes. Although we use additional rendered data, it can achieve training efficiency comparable to ACE. To intuitively compare the localization effects of different methods, we also used NVS to render the images corresponding to the poses estimated by different localization methods, as shown in Figure 6. The upper right is the original image, and the lower left is the image rendered using poses estimated using different methods.
Figure 6: The rendering views obtained based on pose estimations using different localization methods.
Figure 7: Translation error of PoI and C2F with different number of training frames for scence ‘7scenes:fire’.
IV-C Ablation of PoI
To evaluate the effectiveness of our PoI approach, we conducted some experiments in different settings. As shown in Table III, we set the training process using only query data from the training set as the case ‘base’. In this case, the training setting is similar to ACE’s. Case ‘base+poa’ indicates the training with data from the training set and all rendered pixels of the proposed novel pose rendering method. Case ‘base+por’ denotes the pixel-of-random method, which uses randomly sampled pixels from the NVS, and all else are the same as case ‘base+poi’. Case ‘base+poi’ is our final method. The results may show that if we directly use sampled images and the training data without filtering, the results will be worse than the baseline (7scenes). It is easy to understand that because the mapping process is filled with low-quality pixels, which would misguide the network. A concern has been raised that enhanced performance in Novel View Synthesis (NVS) might inversely correlate with the effectiveness of PoI metrics, as improved rendering quality could theoretically reduce the necessity of PoI metrics. We also conduct experiments with different NVS baselines. Case ‘nerf+poi’ represents the implementation of NeRF as the NVS methodology. Experimental results demonstrate that employing 3D Gaussian Splatting (3DGS) for NVS yields superior outcomes compared to NeRF. This observation suggests that the improvement of the effectiveness of the PoI metric will increase with the enhancement of the performance of the NVS method. Experiments also show that images rendered by 3DGS, like those rendered by NeRF-based methods, cannot meet the requirements for direct use by SCR.
IV-D Coarse-to-Fine Experiments with Extremely Sparse Input
To further evaluate the effectiveness of our method, we conduct an additional experiment with sparse input as mentioned in Section III-D. The comparison of translation error between PoI and C2F for scene ‘fire’ is shown in Figure 7. We can see that C2F has an advantage when the input data is sparse.
We use pre-trained MVSplat[42] as the sparse NVS model. To simulate the sparse input, we uniformly resample from the input data every 200 frames. For outdoor datasets like Cambridge Landmarks, we split the data into multiple clusters according to the ground truth translations of camera pose (4 in the experiment) and used only one cluster for training.
The numerical results are shown in Table IV, case ‘base’ denotes the sparse input circumstances with the baseline model. Case ‘coarse’ is the self-pruning step of our method using only rendered data; we still use the Fisher information sampling as the novel pose sampling method. Case ‘c2f’ denotes the fine-tuned results of our proposed method.
TABLE III: Median errors of different implementations of PoI method on 7Scenes and Cambridge dataset.
Method 7Scenes Cambridge Landmarks
[…]
base 1.1cm 0.3° 97.1% 17.7cm 0.3° 32.4%
base+poa 2.3cm 1.58° 89.6% 17.6cm 0.3° 32.2%
base+por 1.4cm 0.46° 95.9% 18.0cm 0.3° 30.4%
nerf+poi 1.1cm 0.4° 95.3% 16.7cm 0.3° 32.3%
base+poi 1.0cm 0.3° 99.1% 15.4cm 0.3° 34.7%
TABLE IV: Median errors of our proposed method with sparse input on 7Scenes and Cambridge dataset.
Method 7Scenes Cambridge Landmarks
[…]
base 2.6cm 0.7° 68.9% 435cm 2.2° 15.7%
coarse 14.3cm 4.4° 14.5% 157cm 1.8° 19.6%
c2f 1.6cm 0.3° 93.0% 22.7cm 0.3° 21.9%
Based on the results, we may find that our fine-tuned model can achieve acceptable results with limited input compared to those that use all the training data.
V Conclusion
In this paper, we propose a PoI method for SCR-based visual localization methods. PoI applies NVS methods, including NeRF and 3DGS, to render synthetic views for training. Experiments show that the localization accuracy improves with the improvement of NVS quality. Furthermore, we designed a coarse-to-fine pipeline for extremely sparse input scenarios. We conducted experiments on indoor and outdoor datasets and achieved state-of-the-art camera pose estimation with comparable training time.